OpenStudy (mathhelp346):

\[\left(\begin{matrix}m\over2\end{matrix}\right)+3=\left(\begin{matrix}m \ \over5\end{matrix}\right)+5\] Explain how you got your answer too @Algebra

OpenStudy (asnaseer):

the smallest number that is a multiple of 5 and 2 is 10, so multiply both sides by 10 to get:\[5m+30=2m+50\]\[5m-2m=50-30\]\[3m=20\]\[m=20/3\]

OpenStudy (mathhelp346):

thanks

OpenStudy (asnaseer):

yw